Find a.b if a=I cap + 2 j cap+ k cap and |b|=3 acting along c vector=icap + j cap+ k cap

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
b = 3 * (i + j + k) / sqrt(1+1+1) = sqrt(3) * (i + j + k)
 
Now
 
a.b = (i + 2j + k) .  sqrt(3) * (i + j + k)
 
 = sqrt(3) * (1 + 2 + 1) = 4 * sqrt(3)